OpenStudy (anonymous):

Can someone help me please... Thanks A new cruise ship line has just launched 3 new ships; the Pacific Paradise, the Carribean Paradise and the Mediterranean Paradise. The Carribean Paradise has 28 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 29 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 881. The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms.

OpenStudy (anonymous):

I would assign variables for the number of rooms for each cruise ship, then change the paragraph into equations.

OpenStudy (anonymous):

The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms.

OpenStudy (noseboy908):

You should assign variables to the ships, not the rooms, since there is only one type of room.

OpenStudy (anonymous):

I am confuse with variables can you help me

OpenStudy (noseboy908):

yeah, in this situation, you have to assign the variables x, y, & z to each (respectably). So, how would you start?

OpenStudy (anonymous):

i have no clue at all.

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

ok, to start out, make x rep. the "Pacific Paradise", y the "Carribean Paradise", & z rep. the "Medditerranean Paradise". You need 3 equations to start, so start simple. we know that all the ships have a total of 881 "Deluxe Rooms", so what do you think could be our first equation?

OpenStudy (anonymous):

x+y+z

OpenStudy (noseboy908):

=?

OpenStudy (anonymous):

881

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):

could you do me a favor and double-check your numbers in the how the problem is written?

OpenStudy (anonymous):

Can someone help me please... Thanks A new cruise ship line has just launched 3 new ships; the Pacific Paradise, the Carribean Paradise and the Mediterranean Paradise. The Carribean Paradise has 28 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 29 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 881. The Pacific Paradise has______deluxe staterooms The Caribbean Paradise has____deluxe The mediterranean Paradise has___deluxe staterooms.

OpenStudy (noseboy908):

yes, but you are certain that the numbers are copied exactly as written in your textbook or workbook?

OpenStudy (anonymous):

Yes.

OpenStudy (noseboy908):

See, I keep coming up with the problem that I never get a whole, non-decimal number whenever I try to solve it. Allow me to explain my process.

OpenStudy (anonymous):

ok

OpenStudy (noseboy908):



OpenStudy (anonymous):

yes

OpenStudy (noseboy908):

Then I used substitution: z-29=4(y+28) Because it's stated in eqn. 1 that x= y+28. distribute: z-29=4y+112 add: z=4y+141. Next, I put (x=y+28), and (z=4y+141) into eqn. 3. To get: (y+28)+y+(4y+141)=881 Add the y's: 6y+169=881 subtract: 6y=712 divide: y=118.66. This is obviously incorrect, since you cannot have a decimal of a room. See any errors?

OpenStudy (anonymous):

no